# algorithmic

### v1

@@ 4.3 Considerations for eigenvector selection in a product-of-manifold model
Consider for example the case for the data illustrated in Figure 4a. Panel (c) shows the leading eight eigenvectors of the graph Laplacian. The leading two eigenvectors are functions of $\theta_{1}$ and by choosing them we completely disregard $\theta_{2}$ with an obvious impact on the feature selection accuracy. A better choice for pseudo-labels would be the first and third eigenvectors, indexed by ${\bm{e}}_{1}$ and ${\bm{e}}_{2}$. Therefore, we need an improved criterion for selecting eigenvectors to serve as pseudolabels for the feature selection process. The following theorem, proven in Appendix A.1, implies that the feature vectors ${\bm{f}}_{i}$ are relevant for developing such a criterion.
@@
In Algorithm 1 we compute the normalized measure of stability for the feature scores $\{s_{m}(h_{i})\}_{m=1}^{p}$ obtained by the model $h_{i}$ to predict the labels computed from the $i$-th eigenvector. When the model $h_{i}$ is linear (or generalized linear), the score is strongly related to the simple inner product of Theorem 2. In that case, Theorem 2 indicates that the inner product between an uniformative eigenvectors and all features is close to zero. Thus, we expect the variance (after normalization) to be similar to the variance of random positive noise. The advantage of the stability measure over the simple linear product as an eigenvector selection criterion is that it allows for more flexibility in the choice of model.
@@ 5.1 Evaluation on real world datasets
We applied SSFS to eight real-world datasets from various domains. Table 4 in Appendix C.1 gives the number of features, samples, and classes in each dataset. All datasets are available online  [^1] https://jundongl.github.io/scikit-feature/datasets.html.
@@
Unlike the evaluation approach taken by Wang et al. [2015], Li et al. [2012], which entailed a grid search over hyper-parameters to report the optimum results for each method, our analysis employed the default hyper-parameters as specified by the respective implementations, including SSFS. This approach aims for a fair comparison to avoid favoring methods that are more sensitive to hyper-parameter adjustments. In addition, it acknowledges the practical constraints in unsupervised settings where hyper-parameter tuning is typically infeasible. Such differences in the approach to hyper-parameter selection could account for discrepancies between the results reported in previous studies and those in our study. See Appendix C for additional details.
@@ Results.
SSFS has been ranked as the best method in three out of eight datasets. It has shown a significant advantage over other competing methods, especially in the Yale, TOX-171, and Prostate-GE datasets. As discussed in Section 3.1, the Prostate-GE dataset has several outliers, and the fourth eigenvector plays a vital role in providing information about the class labels compared to the earlier eigenvectors. SSFS can effectively deal with such challenging scenarios, and this might explain its superior performance. Although our method is not ranked first in the other five datasets, it has produced results comparable to the leading method.
@@ 5.2.1 Synthetic data
We generate a synthetic dataset as follows: the first five features are generated from two isotropic Gaussian blobs; these blobs define the clusters of interest. Additional 45 nuisance features are generated according to a multivariate Gaussian distribution, with zero mean and a block-structured covariance matrix ${\bm{\Sigma}}$, such that each block contains 15 features. The covariance elements ${\bm{\Sigma}}_{i,j}$ are equal to $0.5$ if $i,j$ are in the same block and to $0.01$ otherwise. We generated a total of 500 samples; see Appendix B.1 for further details. In Figure 6a, you can see a scatter plot of the first five features, and in Figure 6b, you can see a visualization of the covariance matrix. Our goal is to identify the features that can distinguish between the two groups.
@@ 5.2.2 Results
We compare to a variation of SSFS termed SSFS (no selection), where we don’t filter the eigenvectors. We train the surrogate feature selector model on the leading $k$ eigenvectors, with $k$ set to the number of distinct classes in the data. Figure 7b, shows that our eigenvector selection scheme provides an advantage in seven out of eight datasets. Similarly to Sec. 5.1, filtering the eigenvectors is especially advantageous on the Prostate-GE dataset, as our method successfully selects the most discriminative eigenvectors (see Figure 2a ). On the synthetic dataset, the selection procedure provides a large advantage, as seen in Table 3. Figure 6c illustrates that the fourth eigenvector is the informative one with respect to the Gaussian blobs. Indeed, the fourth eigenvector and the third eigenvector are selected by the selection procedure. This eigenvector yields better features than MCFS and SSFS (no selection), which rely on the top two eigenvectors.
@@ 6 Discussion and future work
Appendix B Ablation study
B.1 Synthetic data generation
@@
B.2 Detailed Experimental Results
@@
Appendix C Additional Experimental Details
C.1 Datasets
@@
C.2 Hyperparameters

### v2

@@ 4.3 Considerations for eigenvector selection in a product-manifold model
Consider for example the case for the data illustrated in Figure 4a. Panel (b) shows the leading eight eigenvectors of the graph Laplacian. The leading two eigenvectors are functions of $\theta_{1}$ and by choosing them we completely disregard $\theta_{2}$ with an obvious impact on the feature selection accuracy. A better choice for pseudo-labels would be the first and third eigenvectors, indexed by ${\bm{e}}_{1}$ and ${\bm{e}}_{2}$. Therefore, we need an improved criterion for selecting eigenvectors to serve as pseudolabels for the feature selection process. The following theorem, proven in Appendix A.1, implies that the feature vectors ${\bm{f}}_{i}$ are relevant for developing such a criterion.
@@
In Algorithm 1 we compute the normalized measure of stability for the feature scores $\{s_{m}(h_{i})\}_{m=1}^{p}$ obtained by the model $h_{i}$ to predict the labels computed from the $i$-th eigenvector. When the model $h_{i}$ is linear (or generalized linear), the score is strongly related to the simple inner product of Theorem 2. In that case, Theorem 2 indicates that the inner product between uninformative eigenvectors and all features is close to zero. Thus, we expect the variance (after normalization) to be similar to the variance of random positive noise. The advantage of the stability measure over the simple linear product as an eigenvector selection criterion is that it allows for more flexibility in the choice of model.
4.4 Computational Complexity
The computational complexity of our proposed method consists of several key steps. First, constructing the graph representation of the dataset using a Gaussian kernel requires computing pairwise distances between all data points, resulting in a complexity of $O(n^{2}p)$, where $n$ is the number of samples and $p$ is the number of features. Following this, the eigendecomposition of the graph has a complexity of $O(n^{3})$. This can be reduced by randomized methods to $O(n^{2}d)$, since we only require the first $d$ eigenvectors. Once eigenvectors are computed, we apply the $k$-medoids algorithm to cluster the eigenvectors and generate pseudo-labels, adding a further complexity of $O(dkn^{2})$, where $k$ is the number of clusters.
After creating pseudo-labels, we use surrogate models such as logistic regression or gradient-boosted decision trees to evaluate the features. Logistic regression has a complexity of $O(nd)$, while gradient-boosted trees, being more complex, scale with $O(Tn\log n)$, where $T$ is the number of trees. We also use a resampling procedure to estimate the variance of feature importance scores. Each resampling involves training a surrogate model, resulting in an overall complexity of $O(BTn\log n)$ for gradient-boosted trees or $O(Bnd)$ for logistic regression. As a result, the overall computational complexity of SSFS is mainly influenced by the graph construction and training the surrogate mode, resulting in a complexity of $O(n^{2}d+Bnd)$.
@@ 5.1 Evaluation on real world datasets
We applied SSFS to eight real-world datasets from various domains. Table 5 in Appendix F.2 gives the number of features, samples, and classes in each dataset. All datasets are available online  [^1] https://jundongl.github.io/scikit-feature/datasets.html.
@@
Unlike the evaluation approach taken by Wang et al. [2015], Li et al. [2012], which entailed a grid search over hyper-parameters to report the optimum results for each method, our analysis employed the default hyper-parameters as specified by the respective implementations, including SSFS. This approach aims for a fair comparison to avoid favoring methods that are more sensitive to hyper-parameter adjustments. In addition, it acknowledges the practical constraints in unsupervised settings where hyper-parameter tuning is typically infeasible. Such differences in the approach to hyper-parameter selection could account for discrepancies between the results reported in previous studies and those in our study. See Appendix F.1 for additional details.
@@ Results.
SSFS has been ranked as the best method in three out of eight datasets. It has shown a significant advantage over competing methods, especially in the Yale, TOX-171, and Prostate-GE datasets. As discussed in Section 3.1, the Prostate-GE dataset has several outliers, and the fourth eigenvector plays a vital role in providing information about the class labels compared to the earlier eigenvectors. SSFS can effectively deal with such challenging scenarios, and this might explain its superior performance. Although our method is not ranked first in the other five datasets, it has produced results comparable to the leading method.
@@ 5.2.1 Synthetic data
We generate a synthetic dataset as follows: the first five features are generated from two isotropic Gaussian blobs; these blobs define the clusters of interest. Additional 45 nuisance features are generated according to a multivariate Gaussian distribution, with zero mean and a block-structured covariance matrix ${\bm{\Sigma}}$, such that each block contains 15 features. The covariance elements ${\bm{\Sigma}}_{i,j}$ are equal to $0.5$ if $i,j$ are in the same block and $0.01$ otherwise. We generated a total of 500 samples; see Appendix E.1 for further details. In Figure 6a, you can see a scatter plot of the first five features, and in Figure 6b, you can see a visualization of the covariance matrix. Our goal is to identify the features that can distinguish between the two groups.
@@ 5.2.2 Results
We compare to a variation of SSFS termed SSFS (no selection), where we don’t filter the eigenvectors. We train the surrogate feature selector model on the leading $k$ eigenvectors, with $k$ set to the number of distinct classes in the data. Figure 7b shows that our eigenvector selection scheme provides an advantage in seven out of eight datasets. Similarly to Sec. 5.1, filtering the eigenvectors is especially advantageous on the Prostate-GE dataset, as our method successfully selects the most discriminative eigenvectors (see Figure 2a ). On the synthetic dataset, the selection procedure provides a large advantage, as seen in Table 3. Figure 6c illustrates that the fourth eigenvector is the informative one with respect to the Gaussian blobs. Indeed, the fourth eigenvector and the third eigenvector are selected by the selection procedure. This eigenvector yields better features than MCFS and SSFS (no selection), which rely on the top two eigenvectors.
@@ 6 Discussion and future work
Broader Impact Statement
This work involves methods applicable to high-dimensional data, including sensitive domains like healthcare. While our experiments used publicly available datasets, minimizing immediate privacy concerns, practitioners applying this method to sensitive data should ensure compliance with data protection regulations and adopt robust privacy-preserving measures.
There is also a risk of misapplication in critical areas such as medical diagnosis or criminal justice, where biased or incorrect feature selection could lead to harmful decisions. To mitigate this, we recommend integrating bias detection and mitigation techniques into the training pipeline to enhance fairness and reliability.
Careful consideration of these ethical and societal implications is essential for the responsible use of this method.
@@
Appendix B Stability analysis
We perform stability analysis using the Variation of Information (VI) criterion introduced by Meilă [2003], a measure of the distance between two clustering assignments. This criterion can be used to measure the stability of clustering results, which is valuable for assessing feature selection quality. For noisy features, we expect unstable clustering outputs with high variability between runs, while informative features should yield more stable results with less variability.
Our analysis involves applying $k$-means clustering to the selected features 500 times with random initializations, calculating the VI between all pairs of clustering outputs, and analyzing the distribution of VI values. Lower average VI and tighter distribution indicate higher stability, suggesting more informative features. We compared the stability of clusters obtained using the original dataset versus those obtained using features selected by our Stable Spectral Feature Selection (SSFS) method.
Table 4 presents results for the TOX171 and Yale datasets, demonstrating clear improvement in stability when using SSFS. For both datasets, SSFS-selected features yield lower mean VI values, indicating more consistent clustering outcomes. Figure 8 visualizes the distributions of VI values for both datasets, showing the shift towards lower VI values when using SSFS-selected features compared to using all features.
Table 4: Stability Analysis Results: Mean Variation of Information (VI) ± Standard Deviation
Dataset SSFS (top 100 features) All features
TOX171 1.64 ± 0.39 2.01 ± 0.36
Yale 2.12 ± 0.27 2.25 ± 0.27
(a) TOX 171 dataset
(b) Yale dataset
Figure 8: Histograms of VI value distributions for SSFS-selected features (orange) and all features (blue) on (a) TOX 171 and (b) Yale datasets. Lower VI values indicate higher stability in clustering results.
Appendix C Interpretability
Interpretability is a fundamental aspect of unsupervised feature selection, particularly in domains where understanding the meaning and relevance of selected features is crucial.
In the context of our proposed SSFS algorithm, interpretability can be enhanced in several ways. Feature ranking, by examining the weights or importance scores assigned to features during the selection process, can provide insights into which features are most important across the dataset. Visualization techniques, as demonstrated in Figure 9, can reveal patterns and distributions that highlight the discriminative capability of selected features across different classes or clusters. Domain-specific analysis is another effective approach. In biological datasets, for instance, selected features can be mapped back to specific genes or biomarkers, allowing domain experts to validate the biological relevance of the selected features and potentially uncover new insights about the underlying biological processes.
These interpretability techniques enhance the utility of unsupervised feature selection algorithms like SSFS by providing insights into underlying relationships and feature relevance, which are essential for real-world applications across various domains.
Figure 9 visualizes the top twenty features selected by SSFS on the COIL20 dataset. Each column corresponds to a sample, and each row represents a selected feature. The columns are ordered by the class label. The figure highlights the discriminative effect the selected features have over five classes in this dataset.
Figure 9: Visualization of the top 20 selected features by SSFS across five classes in the COIL20 dataset. Samples are grouped by class, separated by white vertical lines, with class labels shown at the top. The color intensity indicates the relative feature values. This visualization allows for the comparison of feature patterns and distributions across different classes, highlighting potential discriminative features and class-specific characteristics.
Appendix D Additional experiments on synthetic data
We conducted further experiments to test our method’s robustness to noise, using two synthetic datasets with 100 samples per trial over 10 trials. Both datasets contain five meaningful features from two isotropic Gaussian blobs but differ in their nuisance features. The first dataset’s nuisance features are sampled from a multivariate Gaussian distribution, similar to Section 5.2.1, but with varying numbers of nuisance features. The second uses a copula of correlated nuisance features formed of marginals with a Laplace distribution. We measure performance using recall, comparing the top five ranked features to the five known meaningful ones.
Results in Figure 10 show SSFS with logistic regression as a linear classifier (denoted by linear_clf) outperforming other methods across both experiments. For Gaussian nuisance features (Figure 10a), SSFS with a linear classifier maintained consistently high performance up to 60 features, while other methods exhibited a more pronounced decline as the number of nuisance features grew. In the case of Laplace-distributed nuisance features (Figure 10b), the advantage of SSFS with a linear classifier was even more evident, with its performance decreasing more slowly as the number of nuisance features increased. Overall, these experiments demonstrate that SSFS is a robust method for identifying relevant features, performing well for various noise distributions and maintaining high recall even with increasing noise levels.
(a) Gaussian nuisance features
(b) Laplace nuisance features
Figure 10: Performance comparison of feature selection methods under increasing noise levels. The x-axis shows the number of nuisance features added to the give meaningful ones, while the y-axis represents the average recall (proportion of the top five ranked features matching the five meaningful features). Each point is the average over 10 trials with 100 samples each. SSFS with a linear classifier (logistic regression) consistently outperforms MCFS and the other SSFS variants.
Appendix E Ablation study additional details
E.1 Synthetic data generation
@@
E.2 Detailed Experimental Results
@@
Appendix F Experiments on real world datasets additional details
F.1 Additional Experimental Details
F.2 Datasets
@@
F.3 Hyperparameters
Hyperparameter selection in unsupervised feature selection presents unique challenges and may be infeasible due to the absence of labeled data for validation. In our methods, we have two types of hyperparameters: (i) For computing the spectral representation (kernel bandwidth and the number of eigenvectors) and (ii) parameters relating to the surrogate model, which predicts the pseudo-labels (i.e., regularization parameter for Logistic Regression). Note that the second type of parameters is used in a prediction task; they can be determined similarly to hyperparameters in supervised settings, for example, through cross-validation.
In our experiments, for the surrogate models, we leverage the default hyperparameters from established machine learning packages such as XGBoost and scikit-learn’s logistic regression, as these are generally robust for a wide range of supervised tasks. However, surrogate models and their hyperparameters can also be chosen by performing hyperparameter tuning with cross-validation based on pseudo-labels derived from the eigenvectors. The selection of surrogate models is flexible and can be additionally guided by domain knowledge. Practitioners can also compare results across various models and analyze their selected features.
The first type of hyperparameters is inherent to the method, for example, the number of considered eigenvectors, which should increase with the number of classes and noise level in the data. For instance, in scenarios with more classes or where noise dominates the spectrum, a larger range of eigenvectors should be considered. While we used 500 resamples in our experiments, this number can be adjusted for larger datasets to balance computational cost and performance. Although traditional hyperparameter tuning may not be feasible in most unsupervised scenarios, practitioners working with domain-specific data could potentially fine-tune parameters using cross-validation on similar, labeled datasets from the same domain (such as images).